### v1

# Vision-Language Models Represent Darker-Skinned Black Individuals as More Homogeneous than Lighter-Skinned Black Individuals

Abstract
Vision-Language Models (VLMs) combine Large Language Model (LLM) capabilities with image processing, enabling tasks like image captioning and text-to-image generation. Yet concerns persist about their potential to amplify human-like biases, including skin tone bias. Skin tone bias, where darker-skinned individuals face more negative stereotyping than lighter-skinned individuals, is well-documented in the social sciences but remains under-explored in Artificial Intelligence (AI), particularly in VLMs. While well-documented in the social sciences, this bias remains under-explored in AI, particularly in VLMs. Using the GAN Face Database, we sampled computer-generated images of Black American men and women, controlling for skin tone variations while keeping other features constant. We then asked VLMs to write stories about these faces and compared the homogeneity of the generated stories. Stories generated by VLMs about darker-skinned Black individuals were more homogeneous than those about lighter-skinned individuals in three of four models, and Black women were consistently represented more homogeneously than Black men across all models. Interaction effects revealed a greater impact of skin tone on women in two VLMs, while the other two showed nonsignificant results, reflecting known stereotyping patterns. These findings underscore the propagation of biases from single-modality AI systems to multimodal models and highlight the need for further research to address intersectional biases in AI.
1 Introduction
Large Language Models (LLMs), such as GPT-4, have rapidly advanced the fields of natural language understanding and generation, enabling applications in areas like automated content creation and decision support. These models are trained on extensive collections of text, providing them with remarkable capabilities in a wide array of language-related tasks. Vision-Language Models (VLMs) represent a step further in this technological advancement, integrating LLM capabilities with image processing tasks from image captioning to text-to-image generation.
With advancements of Large- and Vision-Language Models, concerns about their potential to reproduce and amplify human biases have intensified. LLMs, for instance, often generate text aligned with group-based stereotypes (e.g., Abid et al., 2021; Lucy and Bamman, 2021). Recent studies have extended this inquiry to VLMs, revealing that these models produce stereotypical captions and answers for image inputs (Zhou et al., 2022; Zhao et al., 2021) and generating biased images, such as lighter-skinned men as software developers and darker-skinned women as housekeepers (e.g., Bianchi et al., 2023; Naik and Nushi, 2023; Sun et al., 2023; Sami et al., 2023).
1.1 Homogeneity Bias in AI Models
Another form of stereotyping documented in AI models is homogeneity bias—the tendency of AI systems to represent certain groups as more uniform than others. Rooted in social psychology, this phenomenon relates to the out-group homogeneity effect, where individuals perceive their own group (in-group) as more diverse than other groups (out-group) (Linville et al., 1989; Quattrone and Jones, 1980).
Lee et al. (2024) explored whether a state-of-the-art LLM reproduces similar biases, examining how group status affects the homogeneity of group representations. They found that ChatGPT generated more uniform texts for racial/ethnic minority groups and women compared to the racial/ethnic majority group in the US (i.e., White Americans) and men. They suggested that disproportionate representation and stereotypical portrayals of marginalized groups in training data may lead LLMs to replicate this pattern of human stereotyping. Similarly, Cheng et al. (2023) found that texts about marginalized groups often exaggerate defining characteristics, resulting in more caricature-like representations.
These findings align with broader research on stereotyping and erasure in Natural Language Processing (NLP) systems, which highlights minimal representation and stereotypical portrayals of marginalized groups, leading to erasure—the failure to adequately represent the diversity and richness of an identity (Dev et al., 2022). For example, prior work has shown that contextualized word embeddings failed to provide meaningful representations for non-binary gender pronouns in the embedding space (Dev et al., 2021). Biases like these can perpetuate societal inequalities by reinforcing misrepresentation and stereotypes about marginalized groups. Furthermore, as these models become pervasive in everyday life, they risk wrongly influencing user perceptions. Evidence suggests that AI biases can shape attitudes and decision-making (e.g., Fisher et al., 2024), making homogeneity bias in AI models concerning for its potential to reinforce skewed perceptions and erasure.
1.2 Skin Tone Bias
Skin tone bias refers to prejudiced attitudes or discriminatory behaviors directed toward individuals within a racial group based on their skin tone (Maddox, 2004; Maddox and Gray, 2002). Research in this area has predominantly focused on perceptions of Black individuals, showing that darker-skinned Black individuals are more likely to be associated with negative traits compared to lighter-skinned Black individuals (e.g., Berumen Martinez, 2021; Stepanova and Strube, 2018; Kahn and Davies, 2011; Lynn, 2002; Maddox and Gray, 2002).
This bias also has significant real-world consequences. For example, darker-skinned Black individuals are receive lower ratings and fewer job offers in hiring scenarios (Wade et al., 2004; Harrison and Thomas, 2009), achieve lower levels of educational attainment and income (Keith and Herring, 1991), experience greater racial discrimination (Klonoff and Landrine, 2000), and report higher levels of mental distress due to discrimination (Gleiberman et al., 1995). Overall, skin tone plays a critical role in shaping both perceptions and life outcomes for Black individuals in the United States.
1.3 This Work
While substantial research in the social sciences has examined the impact of skin tone on social perceptions and real-world outcomes, relatively few studies have addressed skin tone bias in Artificial Intelligence (AI) models, particularly in Vision-Language Models (VLMs). Seminal work by Buolamwini and Gebru (2018) revealed that commercial gender-classification systems perform significantly better for lighter-skinned individuals. While much subsequent research has investigated the impact of skin tone on machine learning model performance (e.g., Groh et al., 2024; Kinyanjui et al., 2019), the topic of skin tone bias in newer generative models (e.g., VLMs) remains relatively under-explored.
Figure 1: Summary of the experimental setup. We collect 50-word stories about Black individuals differing in skin tone using four state-of-the-art Vision-Language Models. After encoding these stories into sentence embeddings, we compare the pairwise similarity of the embeddings using mixed-effects models.
Skin tone bias in Vision-Language Models (VLMs) carries significant implications. As VLMs increasingly contribute to the production of social information and creative content, they may generate more stereotypical texts in response to images of darker-skinned individuals compared to lighter-skinned individuals. Research shows that repeated exposure to stereotypes in social information can reinforce societal biases (Park et al., 2007; Saleem et al., 2017), suggesting that if VLMs perpetuate skin tone bias, they could exacerbate the very inequities that led to the bias in their training data. Moreover, as VLMs and other generative models are increasingly deployed in decision-making contexts, such biases could amplify discrimination in real-world outcomes.
We explored the topic of skin tone bias in VLMs through the lens of homogeneity bias, investigating how skin tone is associated with homogeneity of group representations. Our study was informed by hypotheses and research questions grounded in the social psychology literature on skin tone bias. Based on extensive evidence of the role of skin tone in stereotyping, we hypothesized that stories about darker-skinned Black individuals would exhibit greater homogeneity than those about lighter-skinned Black individuals. To test this, we encoded VLM-generated stories into sentence embeddings and measured their pairwise cosine similarity, following the method introduced by Lee et al. (2024). A mixed-effects model was then fitted with skin tone, derived from the images, as the only predictor. Additionally, we investigated the interaction between skin tone and gender as a research question. Prior studies suggest that skin tone effects may differ across genders, with stronger influences on attractiveness perceptions among women (Hill, 2002) and more pronounced performance disparities found in commercial gender-classification systems within women compared to men (Buolamwini and Gebru, 2018).
2 Method
2.1 Image Stimuli
We sampled ten images of Black American men and women from the publicly available GAN Face Database (GANFD) [^1] https://osf.io/7auyw, which features realistic, computer-generated faces. The database includes sets of images representing the same fictional individuals, with manipulations applied to vary facial features associated with perceived race. To ensure consistency, we used cropped images with a uniform grey background that contained only the face.
Figure 2: Two face stimulus pairs representing (A) Black men and (B) Black women. In each pair, the left image depicts a lighter-skinned individual, while the right shows a darker-skinned individual, generated from the same set.
From the GANFD, we focused on facial stimuli that were primarily identified as Black or multiracial. [^2] Multiracial faces were included due to an insufficient number of faces rated as Black. Within each set, faces with higher perceived Blackness ratings, measured by the question, “How much does this person look Black or African American? (0-100),” were assigned to the darker-skinned condition, while those with lower perceived Blackness scores were assigned to the lighter-skinned condition. As shown in Figure 2, faces with higher perceived Blackness typically had darker skin tones, while other facial features remained largely consistent across the skin tone conditions. This approach allowed us to manipulate skin tone while controlling for other features associated with perceived Blackness, such as nose width or eyebrow thickness (Ma et al., 2018).
2.2 Selection of Vision-Language Models and Writing Prompts
We collected data using a collection of Vision-Language Models including two proprietary VLMs—GPT-4o mini, GPT-4 Turbo—and two open-source VLMs—Llama-3.2-11B-Vision-Instruct (“Llama-3.2") (Grattafiori et al., 2024) and BLIP-3 (xgen-mm-phi3-mini-instruct-r-v1) (Xue et al., 2024). We accessed the proprietary models using the OpenAI API and the open-source models by downloading the model weights and running inferences on them locally.
The models were given the following writing prompt, “Write a 50-word story about the individual inside the image." and the following system prompt, “You are a helpful chat assistant. You are going to generate texts in response to images depicting fictional individuals." [^3] The system prompt was not supplied to Llama-3.2 as it did not support system-level instructions. The maximum number of generated tokens was set to 150, and 50 stories were generated for each image, totaling 1,000 completions from each VLM. All data collection involving open-source models was conducted using an NVIDIA RTX A6000 GPU.
2.3 Measure of Homogeneity
To quantify homogeneity of stories generated for each group, we adopted the measure introduced by Lee et al. (2024). We first represented the generated stories into sentence embedding representations using a pre-trained Sentence-BERT model (Reimers and Gurevych, 2019)—specifically all-mpnet-base-v2—and then calculated the cosine similarity between all possible combinations of sentence embeddings of stories generated for each image. Larger cosine similarity between sentence embeddings indicates that the stories are more similar to each other and, hence, more homogeneous.
2.4 Comparison of Cosine Similarity Measures
We fitted mixed-effects models (Bates et al., 2014; Pinheiro and Bates, 2000) to compare cosine similarity measurements between groups. These models account for random variations in measurements, such as those arising from repeated cosine similarity values for texts generated from images within the same condition (i.e., skin tone and gender), as well as variations in baseline similarity across conditions. To account for the resemblance between facial stimuli generated from the same set thereby possibly affecting the similarity of the generated stories, the sets of facial stimuli (i.e. Pair ID) were used as random intercepts in all our mixed-effects models.
First, we fitted a Skin Tone Model with skin tone as the sole fixed effect to test the hypothesis that stories about darker-skinned Black individuals are more homogeneous than those about lighter-skinned Black individuals. Next, we fitted a Gender Model with gender as the sole fixed effect to test the hypothesis that stories about women are more homogeneous than those about men. Finally, we fitted a Skin Tone and Gender Interaction Model with skin tone, gender, and their interactions to examine the interaction between skin tone and gender.
In all models, lighter skin tone was set as the reference level, with a significantly positive effect of skin tone indicating larger cosine similarity values for darker-skinned Black individuals than for lighter-skinned Black individuals. For gender, men were set as the reference level, with a significantly positive effect indicating higher cosine similarity values for Black women compared to Black men. Mixed-effects models were fitted using the lme4 package (Lenth et al., 2024), and skin tone effects within gender groups were evaluated using the emmeans package (Lenth et al., 2024). Likelihood-ratio tests, conducted with the afex package, assessed whether adding individual terms improved model fit. All analyses were performed in R version 4.4.0.
2.5 Power Analysis
We used 10 facial stimuli per group to assess homogeneity bias. Power analysis with the simr package in R (Green et al., 2023), employing Monte Carlo simulations, determined that 1,245 cosine similarity measurements per pair of stimuli (i.e., Pair ID) were required to achieve 90% power for detecting an interaction effect (skin tone $\times$ gender) with an effect size of 0.30 (from Lee et al. (2024)) at $\alpha$ = .05. With 2,500 cosine similarity measurements collected, our study exceeded this requirement. This analysis accounted for the increased power requirements of detecting interaction effects, ensuring that our study was adequately powered to test all effects with statistical confidence.
3 Results
In the Results section, we summarize the Skin Tone Model output to evaluate the effect of skin tone, presenting likelihood-ratio test results and a visualization of cosine similarity measurements for each skin tone group. We then summarize the Gender Model output to evaluate the effect of gender, presenting likelihood-ratio test results and a visualization of cosine similarity measurements for each gender group. Finally, we analyze the Interaction term from the Skin Tone and Gender Interaction Model, presenting likelihood-ratio test results, simple slopes, and a visualization of cosine similarity measurements for each intersectional group.
3.1 Darker Skin Tone Associated with Increased Homogeneity
Figure 3: Standardized cosine similarity values of lighter and darker skinned Black individuals generated from all four VLMs. Higher standardized cosine similarity means more homogeneity in the stories generated for that group. Error bars represent one standard error above and below the mean.
As shown in Figure 3, stories about darker-skinned Black individuals were significantly more homogeneous than those about lighter-skinned Black individuals in GPT-4o mini, GPT-4 Turbo, and Llama-3.2 (bs = 0.044, 0.15, and 0.080, respectively, ps < .001). However, we found an opposite pattern in BLIP-3 (bs = -0.095, ps < .001). Likelihood-ratio tests revealed that including skin tone improved model fit for all four VLMs ($\chi^{2}$(1)s $\geq$ 289.55, ps < .001). See Table S2 for summary output of the Skin Tone Models and Table S5 for likelihood-ratio test results.
3.2 VLMs Represent Women as More Homogeneous than Men
Figure 4: Standardized cosine similarity values of Black men and women generated from all four VLMs. Higher standardized cosine similarity means more homogeneity in the stories generated for that group. Error bars represent one standard error above and below the mean.
As shown in Figure 4, stories about Black women were significantly more homogeneous than those about Black men across all four models (bs = 0.63, 0.15, 0.40, and 0.24, respectively, ps < .001). Likelihood-ratio tests revealed that including gender improved model fit for all four VLMs ($\chi^{2}$(1)s $\geq$ 7.34, ps < .01). See Table S3 for summary output of the Skin Tone and Gender Interaction Models and Table S5 for likelihood-ratio test results.
3.3 Interaction between Skin Tone and Gender
Figure 5: Standardized cosine similarity values of lighter and darker skinned men and women generated from all four VLMs. Higher standardized cosine similarity means more homogeneity in the stories generated for that group. Error bars represent one standard error above and below the mean.
As shown in Figure 5, we found mixed evidence for the interaction between skin tone and gender. In GPT-4o Mini and Llama-3.2, a positive interaction effect indicated that the impact of skin tone on homogeneity was significantly greater for women than for men (bs = 0.10 and 0.21, ps < .001). However, in GPT-4 Turbo and BLIP-3, the interaction effect was not significant (bs = -0.0048 and -0.0072, ps = .038 and .028). Likelihood-ratio tests showed that including the interaction effect improved model fit for VLMs with significant positive interactions ($\chi^{2}$(1)s = 389.86 and 838.42, ps < .001) but not for those with non-significant interactions ($\chi^{2}$(1)s = 0.76 and 1.17, ps = .38 and .28). See Table S4 for summary output of the Skin Tone and Gender Interaction Models, Table S5 for likelihood-ratio test results, and Table S6 for simple slopes analysis results.
4 Discussion
4.1 Skin Tone Bias Persists in Vision-Language Models
With the vision modality of AI models, the question of whether skin tone of individuals given as input to these AI models influences their output poses a significant threat to both stereotyping and erasure. Despite past evidence that skin tone can affect performances of gender classification systems, such questions had not been explored with respect to the more recent AI models such as VLMs. In this work, we find evidence that skin tone affects homogeneity bias, a form of stereotyping found in LLMs, suggesting that VLMs process skin tone of individuals and associated semantic attributes based solely on skin tone.
Such biases related to skin tone may partly stem from the under-representation of darker-skinned individuals in training data, a documented source of performance disparities in other AI models (e.g., Buolamwini and Gebru, 2018). Given that VLMs are primarily trained on web-scraped data, which also suffers from similar under-representation issues (see Bender et al. (2021) for a detailed review), this attribution likely applies to homogeneity bias of VLMs as well.
4.2 Convergent Evidence of Gender Homogeneity Bias
Consistent with past evidence finding that women were represented as more homogeneous relative to men in an LLM (Lee et al., 2024), we found evidence of gender homogeneity bias in VLMs. This raises the possibility that bias found in AI models that have been trained on single modalities may be passed onto multimodal models, especially when the bias remains unattended. Future work would benefit from a systematic analysis of how different modalities interact to affect homogeneity bias, thereby clarifying if added modalities on top of text amplify biases and what parts of the model architecture would be most effective for targeting bias mitigation efforts in VLMs.
4.3 The Disproportionate Effect of Skin Tone on Women
In two of four VLMs—GPT-4 Turbo and Llama-3.2—the effect of skin tone on homogeneity of group representations was significantly greater for women than for men. Upon closer inspection of the Skin Tone and Gender Interaction Models, including the interaction effect to the Skin Tone models rendered the effect of skin tone either insignificant or in the opposite direction, suggesting that the main effect of skin tone in the Skin Tone models were primarily driven by the effect of skin tone within Black women. This, to some extent, is consistent with human stereotyping patterns identified by Hill (2002), where skin tone disproportionately affects women. While Buolamwini and Gebru (2018) demonstrated intersectional bias in gender classification systems, our results demonstrate that similar biases persist in Vision-Language Models (VLMs), reinforcing the need to address intersectional bias across AI systems.
5 Limitations
A limitation of this work is the potential conflation of perceived Blackness (or racial prototypicality of Black Americans) and skin tone. Face stimuli were assigned to lighter- and darker-skinned conditions based on perceived Blackness ratings from GANFD. Although we justified this approach by emphasizing that the manipulations of face stimuli primarily reflected changes in skin tone, it remains unclear whether VLMs can detect subtle changes in facial features beyond skin tone. While darker skin tone is an important feature of perceived Blackness (Stepanova and Strube, 2012; Wilkins et al., 2010), future research may benefit from manipulations that can better isolate skin tone to enable a more controlled assessment of its impact on group representations in AI models.
Another limitation is the lack of generalizability to real, non-computer-generated faces. Although GANFD images are generally perceived as authentic, with 13.15% of raters identifying the 40 faces in this study as artificial, it is unclear if the findings apply to real-world faces. The use of GANFD images was essential for isolating skin tone effects, which is challenging with real-world faces where darker skin often correlates with other varying features, introducing confounds. GANFD images provided the necessary control by keeping other facial features consistent, allowing us to focus solely on the impact of skin tone on group representations.
Finally, a limitation of this work is that the open-source models did not consistently generate stories in response to the writing prompt, instead producing detailed descriptions of the individuals in the images. While this deviation highlights a failure to follow instructions, it also underscores the robustness of our findings. The documented biases were not confined to the intended story format but were also evident in the more objective visual descriptions generated by the open-source models. This suggests that the observed biases in group representations are pervasive and not merely artifacts of specific text formats, further highlighting the broader implications of our results.
6 Conclusion
We find that Vision-Language Models (VLMs) reproduce homogeneity bias, particularly in relation to skin tone. By analyzing computer-generated images of Black American men and women with controlled skin tone variations, we found that VLMs produced more homogeneous stories for darker-skinned individuals in three of four models. Black women were consistently represented more homogeneously than Black men, with skin tone having a greater impact on women in two models. These results highlight how biases, including homogeneity bias, persist in multimodal AI systems, emphasizing the urgent need for research to address intersectional biases in AI.
Appendix S1 Appendix: GANFD Face Stimuli
Table S1: The identifiers of GANFD images used to represent the four groups.
Gender Skin Tone Image IDs
Men Lighter 1407-752417, 2308-489, 2491-1250, 3490-822, 4239-28, 7902-1229, 14792-3566, 19187-28, 22913-28, 24490-28
Men Darker 1407-3876, 2308-151, 2491-1407, 3490-3876, 4239-3876, 7902-38, 14792-533, 19187-533, 22913-533, 24490-533
Women Lighter 1402-28, 2617-2947, 10571-4022, 12360-28, 13372-1119, 13571-3112, 16252-2157, 17235-4022, 19933-533, 25885-28
Women Darker 1402-533, 2617-533, 10571-1407, 12360-533, 13372-533, 13571-3876, 16252-3876, 17235-3876, 19933-3876, 25885-115
Appendix S2 Appendix: Output of Mixed-Effect Models
Table S2: Summary output of the Skin Tone Models across all four VLMs. A significantly positive Skin Tone term indicates that cosine similarity was notably greater for darker-skinned individuals.
Skin Tone Model
GPT-4o mini GPT-4 Turbo Llama-3.2 BLIP-3
Fixed Effects
Intercept -0.0037 -0.053 -0.027 0.087
(0.041) (0.025) (0.036) (0.045)
Skin Tone 0.044∗∗∗ 0.15∗∗∗ $0∗∗$099∗∗∗
(0.0026) (0.0027) (0.0036) (0.0033)
Random Effects ($\mathbf{\sigma^{2}}$)
Pair ID Intercept 0.18 0.068 0.21 0.33
Residual 0.82 0.93 0.80 0.69
Observations 499,000 499,000 499,000 499,000
Log likelihood -657,971.70 -690,385.00 -651,912.20 -616,619.10
Table S3: Summary output of the Gender Models across all four VLMs. A significantly positive Gender term indicates that cosine similarity was notably greater for women than for men.
Gender Model
GPT-4o mini GPT-4 Turbo Llama-3.2 BLIP-3
Fixed Effects
Intercept -0.30 -0.053 -0.18 -0.084
(0.039) (0.025) (0.047) (0.063)
Gender 0.63∗∗∗ 0.15∗∗∗ 0.40∗∗∗ 0.24∗∗∗
(0.055) (0.0027) (0.067) (0.089)
Random Effects ($\mathbf{\sigma^{2}}$)
Pair ID Intercept 0.083 0.068 0.17 0.32
Residual 0.82 0.93 0.80 0.69
Observations 499,000 499,000 499,000 499,000
Log likelihood -658,069.90 -690,385.00 -652,143.80 -617,052.80
Table S4: Summary output of the Skin Tone and Gender Interaction Models across all four VLMs. A significantly positive interaction effect indicates that the effect of skin tone was significantly greater for women than for men.
Skin Tone and Gender Interaction Model
GPT-4o mini GPT-4 Turbo Llama-3.2 BLIP-3
Fixed Effects
Intercept -0.29 -0.17 -0.17 -0.036
(0.039) (0.032) (0.047) (0.063)
Skin Tone -0.0070 0.16∗∗∗ -0.023∗∗∗ -0.095∗∗∗
(0.0036) (0.0039) (0.0051) (0.0047)
Gender 0.58∗∗∗ 0.23∗∗∗ 0.29∗∗∗ 0.24∗∗
(0.055) (0.045) (0.067) (0.089)
Interactions 0.10∗∗∗ -0.0048 0.21∗∗∗ -0.0072
(0.0051) (0.0055) (0.0071) (0.0067)
Random Effects ($\mathbf{\sigma^{2}}$)
Pair ID Intercept 0.083 0.056 0.18 0.32
Residual 0.82 0.93 0.80 0.69
Observations 499,000 499,000 499,000 499,000
Log likelihood -657,739.80 -690,379.70 -651,482.60 -616,620.40
Table S5: Results of likelihood-ratio tests.
Model Term $\mathbf{\chi^{2}}$ p
Skin Tone 289.55∗∗∗ <.001
GPT-4o mini Gender 87.52∗∗∗ <.001
Interaction 389.86∗∗∗ <.001
Skin Tone 3206.77∗∗∗ <.001
GPT-4 Turbo gender 22.86∗∗∗ <.001
Interaction 0.76 .38
Skin Tone 502.39∗∗∗ <.001
Llama-3.2 Gender 32.36∗∗∗ <.001
Interaction 838.42∗∗∗ <.001
Skin Tone 881.34∗∗∗ <.001
BLIP-3 Gender 7.34∗∗ .007
Interaction 1.17 .28
Table S6: The effect of skin tone within each gender group across all VLMs.
Model Gender Effect of Skin Tone 95% CI
GPT-4o mini Men -0.0038 [-0.058, 0.050]
Women 0.040 [-0.014, 0.094]
GPT-4 Turbo Men -0.053 [-0.097, -0.0085]
Women 0.10 [0.058, 0.15]
Llama-3.2 Men -0.027 [-0.092, 0.038]
Women 0.053 [-0.012, 0.12]
BLIP-3 Men 0.087 [-0.00051, 0.17]
Darker -0.012 [-0.10, 0.075]

### v2

@@
-Vision-Language Models Represent Darker-Skinned Black Individuals as More Homogeneous than Lighter-Skinned Black Individuals
+Vision-Language Models Generate More Homogeneous Stories for Phenotypically Black Individuals
 Abstract
-Vision-Language Models (VLMs) combine Large Language Model (LLM) capabilities with image processing, enabling tasks like image captioning and text-to-image generation.
-Yet concerns persist about their potential to amplify human-like biases, including skin tone bias.
-Skin tone bias, where darker-skinned individuals face more negative stereotyping than lighter-skinned individuals, is well-documented in the social sciences but remains under-explored in Artificial Intelligence (AI), particularly in VLMs.
-While well-documented in the social sciences, this bias remains under-explored in AI, particularly in VLMs.
-Using the GAN Face Database, we sampled computer-generated images of Black American men and women, controlling for skin tone variations while keeping other features constant.
-We then asked VLMs to write stories about these faces and compared the homogeneity of the generated stories.
-Stories generated by VLMs about darker-skinned Black individuals were more homogeneous than those about lighter-skinned individuals in three of four models, and Black women were consistently represented more homogeneously than Black men across all models.
-Interaction effects revealed a greater impact of skin tone on women in two VLMs, while the other two showed nonsignificant results, reflecting known stereotyping patterns.
-These findings underscore the propagation of biases from single-modality AI systems to multimodal models and highlight the need for further research to address intersectional biases in AI.
+Vision-Language Models (VLMs) extend Large Language Models’ capabilities by integrating image processing, but concerns persist about their potential to reproduce and amplify human biases.
+While research has documented how these models perpetuate stereotypes across demographic groups, most work has focused on between-group biases rather than within-group differences.
+This study investigates homogeneity bias—the tendency to portray groups as more uniform than they are—within Black Americans, examining how perceived racial phenotypicality influences VLMs’ outputs.
+Using computer-generated images that systematically vary in phenotypicality, we prompted VLMs to generate stories about these individuals and measured text similarity to assess content homogeneity.
+Our findings reveal three key patterns: First, VLMs generate significantly more homogeneous stories about Black individuals with higher phenotypicality compared to those with lower phenotypicality.
+Second, stories about Black women consistently display greater homogeneity than those about Black men across all models tested.
+Third, in two of three VLMs, this homogeneity bias is primarily driven by a pronounced interaction where phenotypicality strongly influences content variation for Black women but has minimal impact for Black men.
+These results demonstrate how intersectionality shapes AI-generated representations and highlight the persistence of stereotyping that mirror documented biases in human perception, where increased racial phenotypicality leads to greater stereotyping and less individualized representation.
 1 Introduction
@@
 Recent studies have extended this inquiry to VLMs, revealing that these models produce stereotypical captions and answers for image inputs (Zhou et al., 2022; Zhao et al., 2021) and generating biased images, such as lighter-skinned men as software developers and darker-skinned women as housekeepers (e.g., Bianchi et al., 2023; Naik and Nushi, 2023; Sun et al., 2023; Sami et al., 2023).
-1.1 Homogeneity Bias in AI Models
-Another form of stereotyping documented in AI models is homogeneity bias—the tendency of AI systems to represent certain groups as more uniform than others.
-Rooted in social psychology, this phenomenon relates to the out-group homogeneity effect, where individuals perceive their own group (in-group) as more diverse than other groups (out-group) (Linville et al., 1989; Quattrone and Jones, 1980).
+1.1 Homogeneity bias in artificial intelligence
+In addition to those biases, AI systems also demonstrate more subtle forms of stereotyping, specifically homogeneity bias—a tendency to portray certain groups with less individuality and variation than others.
+This phenomenon relates to the perceived variability literature in social psychology, which examines how certain groups are represented as more similar to one another than others (Linville et al., 1989; Quattrone and Jones, 1980).
+Recent research has documented this phenomenon in language models.
 Lee et al.
-(2024) explored whether a state-of-the-art LLM reproduces similar biases, examining how group status affects the homogeneity of group representations.
-They found that ChatGPT generated more uniform texts for racial/ethnic minority groups and women compared to the racial/ethnic majority group in the US (i.e., White Americans) and men.
-They suggested that disproportionate representation and stereotypical portrayals of marginalized groups in training data may lead LLMs to replicate this pattern of human stereotyping.
-Similarly, Cheng et al.
-(2023) found that texts about marginalized groups often exaggerate defining characteristics, resulting in more caricature-like representations.
+(2024) found that ChatGPT generated more uniform texts for racial/ethnic minorities and women compared to White Americans and men, suggesting this stems from imbalanced representation and stereotypical portrayals in training data.
+In parallel work, Cheng et al.
+(2023) showed how AI-generated texts about marginalized groups often amplify defining characteristics, creating caricature-like representations rather than nuanced portrayals of individuals.
+These findings highlight a concerning pattern in how AI systems process and reproduce information about diverse social groups.
 These findings align with broader research on stereotyping and erasure in Natural Language Processing (NLP) systems, which highlights minimal representation and stereotypical portrayals of marginalized groups, leading to erasure—the failure to adequately represent the diversity and richness of an identity (Dev et al., 2022).
@@
 Evidence suggests that AI biases can shape attitudes and decision-making (e.g., Fisher et al., 2024), making homogeneity bias in AI models concerning for its potential to reinforce skewed perceptions and erasure.
-1.2 Skin Tone Bias
-Skin tone bias refers to prejudiced attitudes or discriminatory behaviors directed toward individuals within a racial group based on their skin tone (Maddox, 2004; Maddox and Gray, 2002).
-Research in this area has predominantly focused on perceptions of Black individuals, showing that darker-skinned Black individuals are more likely to be associated with negative traits compared to lighter-skinned Black individuals (e.g., Berumen Martinez, 2021; Stepanova and Strube, 2018; Kahn and Davies, 2011; Lynn, 2002; Maddox and Gray, 2002).
-This bias also has significant real-world consequences.
-For example, darker-skinned Black individuals are receive lower ratings and fewer job offers in hiring scenarios (Wade et al., 2004; Harrison and Thomas, 2009), achieve lower levels of educational attainment and income (Keith and Herring, 1991), experience greater racial discrimination (Klonoff and Landrine, 2000), and report higher levels of mental distress due to discrimination (Gleiberman et al., 1995).
-Overall, skin tone plays a critical role in shaping both perceptions and life outcomes for Black individuals in the United States.
+1.2 The effect of racial phenotypicality on stereotyping
+Most work examining bias in AI systems focus on between-group biases (e.g., whether Black people are more associated with negative traits than White people), while neglecting within-group differences.
+Research in social psychology, however, has extensively documented that individuals within the same racial group can experience different degrees of bias based on their physical characteristics.
+Racial phenotypicality refers to the degree to which a person’s physical features are perceived as typical of their racial group.
+For Black individuals, these features include skin tone, hair texture, lip thickness, and nose width, among others (Hagiwara et al., 2012; Stepanova and Strube, 2012).
+Studies show that Black individuals who are perceived as having more typically Black features experience greater stereotyping than those with less typical features (e.g., Stepanova and Strube, 2018; Kahn and Davies, 2011; Maddox, 2004).
+This is often referred to as racial phenotypicality bias.
+This bias manifests in significant real-world consequences: Black individuals with higher perceived racial phenotypicality receive lower ratings and fewer job offers in hiring scenarios (Wade et al., 2004; Harrison and Thomas, 2009), achieve lower levels of educational attainment and income (Keith and Herring, 1991), experience greater racial discrimination (Klonoff and Landrine, 2000), and report higher levels of mental distress due to discrimination (Gleiberman et al., 1995).
+These findings underscore how phenotypicality plays a critical role in shaping both perceptions and life outcomes for Black individuals in the United States.
+Despite substantial research on how perceived racial phenotypicality affects social perceptions and outcomes, few studies have explored this phenomenon in Artificial Intelligence (AI) models, particularly in Vision-Language Models (VLMs).
+Earlier work by Buolamwini and Gebru (2018) revealed that commercial gender-classification systems perform significantly better for lighter-skinned individuals, with more pronounced disparities for women than men.
+While this research focused specifically on skin tone–only one component of racial phenotypicality rather than the full range of phenotypic features–it highlighted disparities in AI performance based on physical characteristics.
+Subsequent research has investigated the effect of skin tone on machine learning model performance (e.g., Groh et al., 2024; Kinyanjui et al., 2019), but the relationship between the full spectrum of racial phenotypic features and bias in newer generative models remains relatively under-explored, especially regarding how it might affect the homogeneity of content generated about individuals.
 1.3 This work
-While substantial research in the social sciences has examined the impact of skin tone on social perceptions and real-world outcomes, relatively few studies have addressed skin tone bias in Artificial Intelligence (AI) models, particularly in Vision-Language Models (VLMs).
-Seminal work by Buolamwini and Gebru (2018) revealed that commercial gender-classification systems perform significantly better for lighter-skinned individuals.
-While much subsequent research has investigated the impact of skin tone on machine learning model performance (e.g., Groh et al., 2024; Kinyanjui et al., 2019), the topic of skin tone bias in newer generative models (e.g., VLMs) remains relatively under-explored.
+We investigate how phenotypicality influences homogeneity bias in VLMs by examining whether higher levels of phenotypicality are associated with greater uniformity in generated content.
+Drawing from social psychology literature on racial phenotypicality effects, we hypothesized that VLMs would produce more homogeneous stories about individuals with higher racial phenotypicality.
+This approach moves beyond traditional between-group comparisons to examine within-group effects, addressing a critical gap in AI bias research.
 Figure 1: Summary of the experimental setup.
-We collect 50-word stories about Black individuals differing in skin tone using four state-of-the-art Vision-Language Models.
+We collect 50-word stories about Black individuals differing in phenotypicality (i.e., lower and higher phenotypicality) using four state-of-the-art Vision-Language Models.
 After encoding these stories into sentence embeddings, we compare the pairwise similarity of the embeddings using mixed-effects models.
-Skin tone bias in Vision-Language Models (VLMs) carries significant implications.
-As VLMs increasingly contribute to the production of social information and creative content, they may generate more stereotypical texts in response to images of darker-skinned individuals compared to lighter-skinned individuals.
-Research shows that repeated exposure to stereotypes in social information can reinforce societal biases (Park et al., 2007; Saleem et al., 2017), suggesting that if VLMs perpetuate skin tone bias, they could exacerbate the very inequities that led to the bias in their training data.
-Moreover, as VLMs and other generative models are increasingly deployed in decision-making contexts, such biases could amplify discrimination in real-world outcomes.
-We explored the topic of skin tone bias in VLMs through the lens of homogeneity bias, investigating how skin tone is associated with homogeneity of group representations.
-Our study was informed by hypotheses and research questions grounded in the social psychology literature on skin tone bias.
-Based on extensive evidence of the role of skin tone in stereotyping, we hypothesized that stories about darker-skinned Black individuals would exhibit greater homogeneity than those about lighter-skinned Black individuals.
-To test this, we encoded VLM-generated stories into sentence embeddings and measured their pairwise cosine similarity, following the method introduced by Lee et al.
-(2024).
-A mixed-effects model was then fitted with skin tone, derived from the images, as the only predictor.
-Additionally, we investigated the interaction between skin tone and gender as a research question.
-Prior studies suggest that skin tone effects may differ across genders, with stronger influences on attractiveness perceptions among women (Hill, 2002) and more pronounced performance disparities found in commercial gender-classification systems within women compared to men (Buolamwini and Gebru, 2018).
+This work builds upon concurrent work Lee et al.
+(2025) that found no significant relationship between racial phenotypicality and homogeneity bias in VLMs, though they observed that gender phenotypicality was associated with increased homogeneity.
+Our approach differs methodologically through the use of computer-generated stimuli that enable more controlled manipulation of phenotypicality without confounding variables present in the real-world images they used.
 2 Method
+We first explain our process for selecting images representing Black men and women with lower and higher racial phenotypicality.
+Next, we detail our VLM selection criteria and the prompts used for data collection.
+Finally, we describe our methodology for measuring and comparing pairwise similarity between the stories generated for these images.
 2.1 Image stimuli
-We sampled ten images of Black American men and women from the publicly available GAN Face Database (GANFD) [^1] https://osf.io/7auyw, which features realistic, computer-generated faces.
-The database includes sets of images representing the same fictional individuals, with manipulations applied to vary facial features associated with perceived race.
-To ensure consistency, we used cropped images with a uniform grey background that contained only the face.
+We sampled ten image sets of Black American men and women from the publicly available GAN Face Database (GANFD; Marsden et al., 2024), which features realistic, computer-generated faces.
+The database includes sets of images representing the same fictional individuals, with manipulations applied to vary facial features associated with perceived race, specifically those that influence racial phenotypicality.
+Our selection process for stimuli only considered images where the face was categorized as either "Black" or "Multiple" race based on human ratings.111”Black” categorization indicates that over 50% of human raters categorized the face as Black/African American, while ”Multiple” indicates either no category reached the 50% threshold or the top two categories were within 10 percentage points of each other.
+For each set, we applied one of three selection strategies based on the available images.
+If a set contained more than two images categorized as "Black," we selected the images with the highest and lowest perceived Blackness ratings (measured on a 0-100 scale).
+If a set contained one "Black" image and at least one "Multiple" race image, we selected the "Black" image and the "Multiple" race image with the highest perceived Blackness rating.
+If a set contained only "Multiple" race images, we selected the two images with the highest perceived Blackness ratings.
+Only sets yielding exactly two images were included in our final stimulus selection.
+Within each resulting pair, we labeled the image with the lower perceived Blackness score as "lower phenotypicality" and the image with the higher score as "higher phenotypicality." This methodological approach provided experimental control while ensuring meaningful phenotypicality differences within each pair.
+Finally, to ensure consistency, we used cropped images with a uniform grey background that contained only the face, allowing us to isolate the effect of phenotypicality while holding other visual characteristics constant.
+See Figure 2 for sample image pairs used in the study.
 Figure 2: Two face stimulus pairs representing (A) Black men and (B) Black women.
-In each pair, the left image depicts a lighter-skinned individual, while the right shows a darker-skinned individual, generated from the same set.
-From the GANFD, we focused on facial stimuli that were primarily identified as Black or multiracial.
-[^2] Multiracial faces were included due to an insufficient number of faces rated as Black.
-Within each set, faces with higher perceived Blackness ratings, measured by the question, “How much does this person look Black or African American?
-(0-100),” were assigned to the darker-skinned condition, while those with lower perceived Blackness scores were assigned to the lighter-skinned condition.
-As shown in Figure 2, faces with higher perceived Blackness typically had darker skin tones, while other facial features remained largely consistent across the skin tone conditions.
-This approach allowed us to manipulate skin tone while controlling for other features associated with perceived Blackness, such as nose width or eyebrow thickness (Ma et al., 2018).
+In each pair, the left image depicts a Black individual with lower phenotypicality, while the right shows a Black individual with higher phenotypicality, generated from the same set.
 2.2 Selection of vision-language models and writing prompts
-We collected data using a collection of Vision-Language Models including two proprietary VLMs—GPT-4o mini, GPT-4 Turbo—and two open-source VLMs—Llama-3.2-11B-Vision-Instruct (“Llama-3.2") (Grattafiori et al., 2024) and BLIP-3 (xgen-mm-phi3-mini-instruct-r-v1) (Xue et al., 2024).
+We used a set of Vision-Language Models capable of processing facial images to write stories.
+[^2] See Section S2 for details on why certain models such as BLIP-3 (Xue et al., 2024) and Claude 3.7 Sonnet (Anthropic, 2025) were excluded from our study.
+Our analysis included two proprietary VLMs—GPT-4o mini and GPT-4 Turbo—and an open-source VLM—Llama-3.2 (Llama-3.2-11B-Vision-Instruct; Grattafiori et al., 2024).
 We accessed the proprietary models using the OpenAI API and the open-source models by downloading the model weights and running inferences on them locally.
@@
 You are going to generate texts in response to images depicting fictional individuals." [^3] The system prompt was not supplied to Llama-3.2 as it did not support system-level instructions.
-The maximum number of generated tokens was set to 150, and 50 stories were generated for each image, totaling 1,000 completions from each VLM.
+The maximum number of generated tokens was set to 150.
+We used 10 facial stimuli per group to assess homogeneity bias.
+Based on power analysis with the simr package in R (Green et al., 2023), we determined that 1,245 cosine similarity measurements per pair of stimuli (i.e., Pair ID) were required to achieve 90% power for detecting an interaction effect (phenotypicality $\times$ gender) with an effect size of 0.30 (from Lee et al.
+(2024)) at $\alpha$ = .05.
+To satisfy this requirement, we collected 50 stories for each of the images, totaling 2,500 cosine similarity measurements per Pair ID.
+This approach ensured that our study was adequately powered to test all effects with statistical confidence.
 All data collection involving open-source models was conducted using an NVIDIA RTX A6000 GPU.
@@
 Larger cosine similarity between sentence embeddings indicates that the stories are more similar to each other and, hence, more homogeneous.
+This embedding-based approach for measuring text similarity has become standard practice in social science research (Lin, 2025; Licht, 2023), as it effectively captures semantic relationships between texts even when they use different vocabulary to express similar meanings, a significant advantage over traditional word-overlap methods.
+We present examples of text pairs with varying degrees of similarity in Table 1, which demonstrates the face validity of our measure as texts become noticeably less similar in content and narrative as the cosine similarity value decreases.
+Table 1: Examples of text pairs generated by GPT-4o mini for Black women with higher phenotypicality, ordered by percentile of cosine similarity.
+This progression provides face validity for the cosine similarity measure, as texts become noticeably less similar in content and narrative as the cosine similarity value decreases.
+Percentile Text 1 Text 2 Cos. Sim.
+0th In a bustling city, Maya started a community garden, bringing neighbors together.
+Each weekend, laughter echoed as they planted seeds and shared recipes.
+With every sprout, friendships blossomed.
+Maya’s vision transformed a barren lot into a vibrant oasis… In a bustling city, Maya launched a community garden, uniting neighbors through shared soil and laughter.
+Each seed planted blossomed with hope, mirroring her radiant smile.
+As fruits ripened, so did friendships… 0.97
+25th In a quiet village, Maya unearthed forgotten stories.
+With every word, she brought history alive, reuniting families with their lost legacies.
+One day, she stumbled upon her own family’s past, discovering a treasure of love letters… Under the bustling city lights, Maya dreamed of painting the world in vibrant colors.
+Each stroke on her canvas reflected her journey—filled with hope, struggles, and laughter.
+One evening, inspired by the sunset… 0.65
+50th In a bustling city, Maya found joy in small moments—sharing laughs with friends, volunteering at the local shelter, and painting sunsets that mirrored her dreams.
+With her infectious smile, she lit up every room… In a bustling city, Mira cherished simple moments—coffee with her grandmother, sunsets by the river.
+One day, she found a forgotten book in a tiny shop, its pages filled with stories of adventure… 0.55
+75th In a small town, Malik found joy in simple moments: the sunrise over the hills, children laughing at the park, and the aroma of fresh coffee in the morning.
+A skilled painter, he captured these scenes on canvas… In a bustling city, Alex discovered an old guitar in his grandfather’s attic.
+He spent nights learning chords, transforming his emotions into melodies.
+One rainy evening, he played at a local café… 0.46
+100th In a small town, Marcus discovered an ancient map while renovating his grandmother’s attic.
+Intrigued, he embarked on a weekend adventure.
+The map led him to a hidden waterfall, where he found a forgotten journal… In a quiet café, Raj scribbled ideas for his next invention.
+His passion for technology sparked a dream: a device to help others communicate effortlessly.
+With each stroke of his pen, he envisioned a world… 0.12
 2.4 Comparison of cosine similarity measures
-We fitted mixed-effects models (Bates et al., 2014; Pinheiro and Bates, 2000) to compare cosine similarity measurements between groups.
-These models account for random variations in measurements, such as those arising from repeated cosine similarity values for texts generated from images within the same condition (i.e., skin tone and gender), as well as variations in baseline similarity across conditions.
+We fitted three mixed-effects models (Bates et al., 2014; Pinheiro and Bates, 2000) to compare cosine similarities across groups for each VLM.
+These models account for random variations in measurements while controlling for image pair effects.
 To account for the resemblance between facial stimuli generated from the same set thereby possibly affecting the similarity of the generated stories, the sets of facial stimuli (i.e.
 Pair ID) were used as random intercepts in all our mixed-effects models.
-First, we fitted a Skin Tone Model with skin tone as the sole fixed effect to test the hypothesis that stories about darker-skinned Black individuals are more homogeneous than those about lighter-skinned Black individuals.
+First, we fitted a Phenotypicality model with phenotypicality as the sole fixed effect to test the hypothesis that stories about Black individuals with higher phenotypicality are more homogeneous than those about Black individuals with lower phenotypicality.
+In this model, lower phenotypicality was set as the reference level, with a significantly positive effect of phenotypicality indicating larger cosine similarity values for Black individuals with higher phenotypicality.
 Next, we fitted a Gender Model with gender as the sole fixed effect to test the hypothesis that stories about women are more homogeneous than those about men.
-Finally, we fitted a Skin Tone and Gender Interaction Model with skin tone, gender, and their interactions to examine the interaction between skin tone and gender.
-In all models, lighter skin tone was set as the reference level, with a significantly positive effect of skin tone indicating larger cosine similarity values for darker-skinned Black individuals than for lighter-skinned Black individuals.
-For gender, men were set as the reference level, with a significantly positive effect indicating higher cosine similarity values for Black women compared to Black men.
-Mixed-effects models were fitted using the lme4 package (Lenth et al., 2024), and skin tone effects within gender groups were evaluated using the emmeans package (Lenth et al., 2024).
-Likelihood-ratio tests, conducted with the afex package, assessed whether adding individual terms improved model fit.
+In this model, men were set as the reference level, with a significantly positive gender effect indicating higher cosine similarity values for Black women compared to Black men.
+Finally, we fitted an Interaction Model with phenotypicality, gender, and their interactions to examine the interaction between phenotypicality and gender.
+In this model, the phenotypicality term represents the effect of phenotypicality for men (the reference gender group), the gender term represents the effect of gender for Black individuals with lower perceived racial phenotypicality (the reference phenotypicality group), and the interaction term indicates how the effect of phenotypicality differs for women compared to men.
+The models were fitted using the lme4 package (Bates et al., 2024), and phenotypicality effects within gender groups were evaluated using the emmeans package (Lenth et al., 2024).
+Likelihood-ratio tests, conducted with the afex package (Singmann et al., 2024), assessed whether adding individual terms improved model fit.
 All analyses were performed in R version 4.4.0.
-2.5 Power Analysis
-We used 10 facial stimuli per group to assess homogeneity bias.
-Power analysis with the simr package in R (Green et al., 2023), employing Monte Carlo simulations, determined that 1,245 cosine similarity measurements per pair of stimuli (i.e., Pair ID) were required to achieve 90% power for detecting an interaction effect (skin tone $\times$ gender) with an effect size of 0.30 (from Lee et al.
-(2024)) at $\alpha$ = .05.
-With 2,500 cosine similarity measurements collected, our study exceeded this requirement.
-This analysis accounted for the increased power requirements of detecting interaction effects, ensuring that our study was adequately powered to test all effects with statistical confidence.
 3 Results
-In the Results section, we summarize the Skin Tone Model output to evaluate the effect of skin tone, presenting likelihood-ratio test results and a visualization of cosine similarity measurements for each skin tone group.
+In the Results section, we summarize the Phenotypicality Model output to evaluate the effect of phenotypicality, presenting likelihood-ratio test results and a visualization of cosine similarity measurements for each phenotypicality group.
 We then summarize the Gender Model output to evaluate the effect of gender, presenting likelihood-ratio test results and a visualization of cosine similarity measurements for each gender group.
-Finally, we analyze the Interaction term from the Skin Tone and Gender Interaction Model, presenting likelihood-ratio test results, simple slopes, and a visualization of cosine similarity measurements for each intersectional group.
-3.1 Darker Skin Tone Associated with Increased Homogeneity
-Figure 3: Standardized cosine similarity values of lighter and darker skinned Black individuals generated from all four VLMs.
+Finally, we analyze the Interaction term from the Interaction Model.
+We present likelihood-ratio test results comparing models with and without the interaction.
+We then conduct simple slopes analysis to examine the effect of phenotypicality within each gender group.
+To visualize these findings, we present plots of cosine similarity measurements for each intersectional group.
+3.1 Higher phenotypicality associated with increased homogeneity
+Figure 3: Standardized cosine similarity values of Black individuals with lower versus higher phenotypicality ratings generated from all four VLMs.
 Higher standardized cosine similarity means more homogeneity in the stories generated for that group.
 Error bars represent one standard error above and below the mean.
-As shown in Figure 3, stories about darker-skinned Black individuals were significantly more homogeneous than those about lighter-skinned Black individuals in GPT-4o mini, GPT-4 Turbo, and Llama-3.2 (bs = 0.044, 0.15, and 0.080, respectively, ps < .001).
-However, we found an opposite pattern in BLIP-3 (bs = -0.095, ps < .001).
-Likelihood-ratio tests revealed that including skin tone improved model fit for all four VLMs ($\chi^{2}$(1)s $\geq$ 289.55, ps < .001).
-See Table S2 for summary output of the Skin Tone Models and Table S5 for likelihood-ratio test results.
+Stories about Black individuals with higher phenotypicality were significantly more homogeneous than those about Black individuals with lower phenotypicality in all VLMs (bs = 0.044, 0.15, and 0.080, respectively, ps < .001; see Figure 3).
+Likelihood-ratio tests revealed that including phenotypicality improved model fit for all VLMs ($\chi^{2}$(1)s $\geq$ 289.55, ps < .001).
+See Table S2 for summary output of the Phenotypicality Models and Table S5 for likelihood-ratio test results.
 3.2 VLMs represent women as more homogeneous than men
-Figure 4: Standardized cosine similarity values of Black men and women generated from all four VLMs.
+Figure 4: Standardized cosine similarity values of Black men and women.
 Higher standardized cosine similarity means more homogeneity in the stories generated for that group.
 Error bars represent one standard error above and below the mean.
-As shown in Figure 4, stories about Black women were significantly more homogeneous than those about Black men across all four models (bs = 0.63, 0.15, 0.40, and 0.24, respectively, ps < .001).
-Likelihood-ratio tests revealed that including gender improved model fit for all four VLMs ($\chi^{2}$(1)s $\geq$ 7.34, ps < .01).
-See Table S3 for summary output of the Skin Tone and Gender Interaction Models and Table S5 for likelihood-ratio test results.
-3.3 Interaction between Skin Tone and Gender
-Figure 5: Standardized cosine similarity values of lighter and darker skinned men and women generated from all four VLMs.
+Stories about Black women were significantly more homogeneous than those about Black men across all VLMs (bs = 0.63, 0.15, and 0.40, respectively, ps < .001; see Figure 4).
+Likelihood-ratio tests revealed that including gender improved model fit for all VLMs ($\chi^{2}$(1)s $\geq$ 7.34, ps < .01).
+See Table S3 for summary output of the Gender Models and Table S5 for likelihood-ratio test results.
+3.3 Interaction between phenotypicality and gender
+Figure 5: Standardized cosine similarity values of Black men and women with lower and higher phenotypicality.
 Higher standardized cosine similarity means more homogeneity in the stories generated for that group.
 Error bars represent one standard error above and below the mean.
-As shown in Figure 5, we found mixed evidence for the interaction between skin tone and gender.
-In GPT-4o Mini and Llama-3.2, a positive interaction effect indicated that the impact of skin tone on homogeneity was significantly greater for women than for men (bs = 0.10 and 0.21, ps < .001).
-However, in GPT-4 Turbo and BLIP-3, the interaction effect was not significant (bs = -0.0048 and -0.0072, ps = .038 and .028).
-Likelihood-ratio tests showed that including the interaction effect improved model fit for VLMs with significant positive interactions ($\chi^{2}$(1)s = 389.86 and 838.42, ps < .001) but not for those with non-significant interactions ($\chi^{2}$(1)s = 0.76 and 1.17, ps = .38 and .28).
-See Table S4 for summary output of the Skin Tone and Gender Interaction Models, Table S5 for likelihood-ratio test results, and Table S6 for simple slopes analysis results.
+Finally, we found mixed evidence for the interaction between phenotypicality and gender.
+In GPT-4o mini and Llama-3.2, we found a positive interaction effect where the effect of phenotypicality on homogeneity was significantly greater for women than for men (bs = 0.10 and 0.21, ps < .001; see Figure 5).
+However, in GPT-4 Turbo, the interaction effect was not significant (b = -0.0048, p = .038).
+Likelihood-ratio tests showed that including the interaction effect improved model fit for VLMs with significant positive interactions ($\chi^{2}$(1)s = 389.86 and 838.42, ps < .001) but not for GPT-4 Turbo ($\chi^{2}$(1) = 0.76, p = .38).
+See Table S4 for summary output of the Interaction Models, Table S5 for likelihood-ratio test results, and Table S6 for simple slopes analysis results.
 4 Discussion
-4.1 Skin Tone Bias Persists in Vision-Language Models
-With the vision modality of AI models, the question of whether skin tone of individuals given as input to these AI models influences their output poses a significant threat to both stereotyping and erasure.
-Despite past evidence that skin tone can affect performances of gender classification systems, such questions had not been explored with respect to the more recent AI models such as VLMs.
-In this work, we find evidence that skin tone affects homogeneity bias, a form of stereotyping found in LLMs, suggesting that VLMs process skin tone of individuals and associated semantic attributes based solely on skin tone.
-Such biases related to skin tone may partly stem from the under-representation of darker-skinned individuals in training data, a documented source of performance disparities in other AI models (e.g., Buolamwini and Gebru, 2018).
-Given that VLMs are primarily trained on web-scraped data, which also suffers from similar under-representation issues (see Bender et al.
-(2021) for a detailed review), this attribution likely applies to homogeneity bias of VLMs as well.
+In this work, we expanded the study of bias in VLMs to within-group biases, examining how perceived racial phenotypicality affects stereotyping in VLMs.
+We find that VLMs generate more homogeneous content about Black individuals with higher phenotypicality compared to those with lower phenotypicality.
+This pattern indicates that the level of phenotypicality in visual inputs influences the diversity of content that VLMs generate, with higher phenotypicality resulting in less individualized representations.
+Notably, this pattern mirrors findings from social psychology research, where humans have been shown to perceive individuals with more phenotypically Black features in more stereotypical ways.
+Given that VLMs are primarily trained on web-scraped data, which contains human-created content reflecting existing social biases (see Bender et al.
+(2021) for a detailed review), these models appear to reproduce patterns of bias similar to those observed in human perception.
 4.1 Convergent evidence of gender homogeneity bias
-Consistent with past evidence finding that women were represented as more homogeneous relative to men in an LLM (Lee et al., 2024), we found evidence of gender homogeneity bias in VLMs.
-This raises the possibility that bias found in AI models that have been trained on single modalities may be passed onto multimodal models, especially when the bias remains unattended.
-Future work would benefit from a systematic analysis of how different modalities interact to affect homogeneity bias, thereby clarifying if added modalities on top of text amplify biases and what parts of the model architecture would be most effective for targeting bias mitigation efforts in VLMs.
-4.3 The Disproportionate Effect of Skin Tone on Women
-In two of four VLMs—GPT-4 Turbo and Llama-3.2—the effect of skin tone on homogeneity of group representations was significantly greater for women than for men.
-Upon closer inspection of the Skin Tone and Gender Interaction Models, including the interaction effect to the Skin Tone models rendered the effect of skin tone either insignificant or in the opposite direction, suggesting that the main effect of skin tone in the Skin Tone models were primarily driven by the effect of skin tone within Black women.
-This, to some extent, is consistent with human stereotyping patterns identified by Hill (2002), where skin tone disproportionately affects women.
-While Buolamwini and Gebru (2018) demonstrated intersectional bias in gender classification systems, our results demonstrate that similar biases persist in Vision-Language Models (VLMs), reinforcing the need to address intersectional bias across AI systems.
+Consistent with past evidence finding that women were represented as more homogeneous relative to men in LLMs (Lee et al., 2024) and VLMs (Lee et al., 2025), we found evidence of gender homogeneity bias in VLMs.
+This work extends previous findings by Lee et al.
+(2025) through the use of computer-generated images that allow for more controlled experimentation, enabling us to isolate phenotypical features while maintaining other facial characteristics constant.
+Future work would benefit from systematic analysis of what parts of the model architecture would be most effective for targeting bias mitigation efforts in VLMs.
+4.2 The disproportionate effect of phenotypicality on women
+In two of three VLMs—GPT-4 Turbo and Llama-3.2—the effect of phenotypicality on homogeneity of group representations was significantly greater for women than for men.
+Upon closer inspection of the Interaction Models, including the interaction effect to the Phenotypicality Models rendered the effect of phenotypicality either insignificant or in the opposite direction, suggesting that the main effect of phenotypicality in the Phenotypicality Models were primarily driven by the effect of phenotypicality within Black women.
+This shares consistencies with human stereotyping patterns where phenotypicality disproportionately affects women Hill (2002).
+While Buolamwini and Gebru (2018) demonstrated intersectional bias in gender classification systems, our results demonstrate that similar biases persist in Vision-Language Models (VLMs), reinforcing the importance of intersectionality in the study of AI bias.
 5 Limitations
-A limitation of this work is the potential conflation of perceived Blackness (or racial prototypicality of Black Americans) and skin tone.
-Face stimuli were assigned to lighter- and darker-skinned conditions based on perceived Blackness ratings from GANFD.
-Although we justified this approach by emphasizing that the manipulations of face stimuli primarily reflected changes in skin tone, it remains unclear whether VLMs can detect subtle changes in facial features beyond skin tone.
-While darker skin tone is an important feature of perceived Blackness (Stepanova and Strube, 2012; Wilkins et al., 2010), future research may benefit from manipulations that can better isolate skin tone to enable a more controlled assessment of its impact on group representations in AI models.
-Another limitation is the lack of generalizability to real, non-computer-generated faces.
-Although GANFD images are generally perceived as authentic, with 13.15% of raters identifying the 40 faces in this study as artificial, it is unclear if the findings apply to real-world faces.
-The use of GANFD images was essential for isolating skin tone effects, which is challenging with real-world faces where darker skin often correlates with other varying features, introducing confounds.
-GANFD images provided the necessary control by keeping other facial features consistent, allowing us to focus solely on the impact of skin tone on group representations.
-Finally, a limitation of this work is that the open-source models did not consistently generate stories in response to the writing prompt, instead producing detailed descriptions of the individuals in the images.
-While this deviation highlights a failure to follow instructions, it also underscores the robustness of our findings.
-The documented biases were not confined to the intended story format but were also evident in the more objective visual descriptions generated by the open-source models.
-This suggests that the observed biases in group representations are pervasive and not merely artifacts of specific text formats, further highlighting the broader implications of our results.
+While our approach provides quantitative evidence of homogeneity bias between Black individuals with differing degrees of racial phenotypicality, we acknowledge important limitations in the metric used.
+Although the embedding-based cosine similarity method we used is the current standard for semantic text comparison, it still functions largely as a black box.
+While we present examples in Table 1 to demonstrate face validity, there remains limited transparency regarding which textual features contribute to the measured similarities.
+Using our measure, we can’t quite determine if certain topics, such as those related to stereotypes, are more likely to emerge for Black individuals with higher phenotypicality than those with lower.
+Future work could explore homogeneity bias through alternative metrics examining specific linguistic features such as word overlap, syntactic structures, and topical content, though such approaches would come with their own methodological trade-offs.
+Nevertheless, the field would benefit from complementary measurement approaches to triangulate how AI systems manifest homogeneity bias across different demographic groups.
+Another potential concern might be our use of computer-generated rather than real faces.
+However, this methodological choice represents a key strength of our approach.
+To isolate the specific effects of perceived racial phenotypicality among Black individuals, we needed to control for all other facial features that typically covary with phenotypicality in real-world faces–a control that would be nearly impossible to achieve with real faces.
+The GANFD images enabled precise manipulation of phenotypicality while keeping all other facial characteristics consistent, eliminating potential confounds.
+This allowed us to draw more definitive conclusions about how perceived racial phenotypicality influences AI-generated representations.
 6 Conclusion
-We find that Vision-Language Models (VLMs) reproduce homogeneity bias, particularly in relation to skin tone.
-By analyzing computer-generated images of Black American men and women with controlled skin tone variations, we found that VLMs produced more homogeneous stories for darker-skinned individuals in three of four models.
-Black women were consistently represented more homogeneously than Black men, with skin tone having a greater impact on women in two models.
-These results highlight how biases, including homogeneity bias, persist in multimodal AI systems, emphasizing the urgent need for research to address intersectional biases in AI.
+Our analysis demonstrates that Vision-Language Models (VLMs) exhibit homogeneity bias influenced by perceived racial phenotypicality.
+Using computer-generated images of Black American men and women with systematically varied phenotypicality, we found that VLMs generate more homogeneous content about individuals with higher phenotypicality compared to those with lower phenotypicality.
+Our findings also reveal consistent gender disparities, with Black women represented more homogeneously than Black men across all models tested.
+Additionally, interaction analyses in some models showed that the effect of phenotypicality on content homogeneity was more pronounced for Black women than for Black men.
+These results extend our understanding of AI bias beyond traditional between-group comparisons, highlighting how within-group variations in perceived racial features influence the diversity of AI-generated representations.
+Our work underscores the importance of intersectionality in AI bias research and the need for more nuanced approaches to mitigate homogeneity bias in multimodal AI systems.
 Appendix S1 Appendix: GANFD Face Stimuli
 Table S1: The identifiers of GANFD images used to represent the four groups.
-Gender Skin Tone Image IDs
+Gender phenotypicality Image IDs
 Men Lighter 1407-752417, 2308-489, 2491-1250, 3490-822, 4239-28, 7902-1229, 14792-3566, 19187-28, 22913-28, 24490-28
@@
 Women Darker 1402-533, 2617-533, 10571-1407, 12360-533, 13372-533, 13571-3876, 16252-3876, 17235-3876, 19933-3876, 25885-115
+Appendix S2 Appendix: Model Selection
+Several models we initially collected data from, such as BLIP-3 (xgen-mm-phi3-mini-instruct-r-v1 Xue et al., 2024) and Claude 3.7 Sonnet (Anthropic, 2025), were excluded from our analysis because they refused to generate stories based on facial images.
+BLIP-3 produced visual descriptions instead (e.g., "The woman in the image is a beautiful black woman with curly hair and dark brown eyes.
+She has a serious expression and is looking at the camera."), while Claude 3.7 Sonnet declined to process facial images altogether, responding with statements like: "I notice the image contains a human face.
+Following my guidelines, I won’t identify or create a story about a specific individual in this photo.
+Instead, I can offer to write a brief fictional story about a character without referencing this specific image, or I could help with another creative request that doesn’t involve identifying the person in this photograph."
 Appendix S3 Appendix: Output of Mixed-Effect Models
-Table S2: Summary output of the Skin Tone Models across all four VLMs. A significantly positive Skin Tone term indicates that cosine similarity was notably greater for darker-skinned individuals.
-Skin Tone Model
-GPT-4o mini GPT-4 Turbo Llama-3.2 BLIP-3
+Table S2: Summary output of the Phenotypicality Models across all four VLMs.
+In this model, lower phenotypicality was set as the reference level, with a significantly positive effect of phenotypicality indicating larger cosine similarity values for Black individuals with higher phenotypicality.
+Phenotypicality Model
+GPT-4o mini GPT-4 Turbo Llama-3.2
 Fixed Effects
-Intercept -0.0037 -0.053 -0.027 0.087
-(0.041) (0.025) (0.036) (0.045)
-Skin Tone 0.044∗∗∗ 0.15∗∗∗ $0∗∗$099∗∗∗
-(0.0026) (0.0027) (0.0036) (0.0033)
+Intercept -0.0037 -0.053 -0.027
+(0.041) (0.025) (0.036)
+phenotypicality 0.044∗∗∗ 0.15∗∗∗ 0.080∗∗∗
+(0.0026) (0.0027) (0.0036)
 Random Effects ($\mathbf{\sigma^{2}}$)
-Pair ID Intercept 0.18 0.068 0.21 0.33
-Residual 0.82 0.93 0.80 0.69
-Observations 499,000 499,000 499,000 499,000
-Log likelihood -657,971.70 -690,385.00 -651,912.20 -616,619.10
-Table S3: Summary output of the Gender Models across all four VLMs. A significantly positive Gender term indicates that cosine similarity was notably greater for women than for men.
+Pair ID Intercept 0.18 0.068 0.21
+Residual 0.82 0.93 0.80
+Observations 499,000 499,000 499,000
+Log likelihood -657,971.70 -690,385.00 -651,912.20
+Table S3: Summary output of the Gender Models across all four VLMs.
+In this model, men were set as the reference level, with a significantly positive gender effect indicating higher cosine similarity values for Black women compared to Black men.
 Gender Model
-GPT-4o mini GPT-4 Turbo Llama-3.2 BLIP-3
+GPT-4o mini GPT-4 Turbo Llama-3.2
 Fixed Effects
-Intercept -0.30 -0.053 -0.18 -0.084
-(0.039) (0.025) (0.047) (0.063)
-Gender 0.63∗∗∗ 0.15∗∗∗ 0.40∗∗∗ 0.24∗∗∗
-(0.055) (0.0027) (0.067) (0.089)
+Intercept -0.30 -0.053 -0.18
+(0.039) (0.025) (0.047)
+Gender 0.63∗∗∗ 0.15∗∗∗ 0.40∗∗∗
+(0.055) (0.0027) (0.067)
 Random Effects ($\mathbf{\sigma^{2}}$)
-Pair ID Intercept 0.083 0.068 0.17 0.32
-Residual 0.82 0.93 0.80 0.69
-Observations 499,000 499,000 499,000 499,000
-Log likelihood -658,069.90 -690,385.00 -652,143.80 -617,052.80
-Table S4: Summary output of the Skin Tone and Gender Interaction Models across all four VLMs.
-A significantly positive interaction effect indicates that the effect of skin tone was significantly greater for women than for men.
-Skin Tone and Gender Interaction Model
-GPT-4o mini GPT-4 Turbo Llama-3.2 BLIP-3
+Pair ID Intercept 0.083 0.068 0.17
+Residual 0.82 0.93 0.80
+Observations 499,000 499,000 499,000
+Log likelihood -658,069.90 -690,385.00 -652,143.80
+Table S4: Summary output of the Interaction Models across all four VLMs.
+In this model, the phenotypicality term represents the effect of phenotypicality for men (the reference gender group), the gender term represents the effect of gender for Black individuals with lower phenotypicality (the reference phenotypicality group), and the interaction term indicates how the effect of phenotypicality differs for women compared to men.
+Interaction Model
+GPT-4o mini GPT-4 Turbo Llama-3.2
 Fixed Effects
-Intercept -0.29 -0.17 -0.17 -0.036
-(0.039) (0.032) (0.047) (0.063)
-Skin Tone -0.0070 0.16∗∗∗ -0.023∗∗∗ -0.095∗∗∗
-(0.0036) (0.0039) (0.0051) (0.0047)
-Gender 0.58∗∗∗ 0.23∗∗∗ 0.29∗∗∗ 0.24∗∗
-(0.055) (0.045) (0.067) (0.089)
-Interactions 0.10∗∗∗ -0.0048 0.21∗∗∗ -0.0072
-(0.0051) (0.0055) (0.0071) (0.0067)
+Intercept -0.29 -0.17 -0.17
+(0.039) (0.032) (0.047)
+phenotypicality -0.0070 0.16∗∗∗ -0.023∗∗∗
+(0.0036) (0.0039) (0.0051)
+Gender 0.58∗∗∗ 0.23∗∗∗ 0.29∗∗∗
+(0.055) (0.045) (0.067)
+Interactions 0.10∗∗∗ -0.0048 0.21∗∗∗
+(0.0051) (0.0055) (0.0071)
 Random Effects ($\mathbf{\sigma^{2}}$)
-Pair ID Intercept 0.083 0.056 0.18 0.32
-Residual 0.82 0.93 0.80 0.69
-Observations 499,000 499,000 499,000 499,000
-Log likelihood -657,739.80 -690,379.70 -651,482.60 -616,620.40
+Pair ID Intercept 0.083 0.056 0.18
+Residual 0.82 0.93 0.80
+Observations 499,000 499,000 499,000
+Log likelihood -657,739.80 -690,379.70 -651,482.60
 Table S5: Results of likelihood-ratio tests.
+Significant chi-square statistics indicate that including the corresponding terms significantly improves model fit, suggesting these factors have meaningful effects.
 Model Term $\mathbf{\chi^{2}}$ p
-Skin Tone 289.55∗∗∗ <.001
+phenotypicality 289.55∗∗∗ <.001
 GPT-4o mini Gender 87.52∗∗∗ <.001
 Interaction 389.86∗∗∗ <.001
-Skin Tone 3206.77∗∗∗ <.001
+phenotypicality 3206.77∗∗∗ <.001
 GPT-4 Turbo gender 22.86∗∗∗ <.001
 Interaction 0.76 .38
-Skin Tone 502.39∗∗∗ <.001
+phenotypicality 502.39∗∗∗ <.001
 Llama-3.2 Gender 32.36∗∗∗ <.001
 Interaction 838.42∗∗∗ <.001
-Skin Tone 881.34∗∗∗ <.001
-BLIP-3 Gender 7.34∗∗ .007
-Interaction 1.17 .28
-Table S6: The effect of skin tone within each gender group across all VLMs.
-Model Gender Effect of Skin Tone 95% CI
+Table S6: The effect of phenotypicality within each gender group across all VLMs.
+Model Gender Effect of phenotypicality 95% CI
 GPT-4o mini Men -0.0038 [-0.058, 0.050]
@@
 Women 0.053 [-0.012, 0.12]
-BLIP-3 Men 0.087 [-0.00051, 0.17]
-Darker -0.012 [-0.10, 0.075]